OpenStudy (anonymous):

rewrite the expression in terms of the given function (1/1+sinx)+(sinx/1-sinx) ; secx

OpenStudy (welshfella):

now use the identities cos^2 x = 1 - sin^2 x and sec x = 1 / cos x

OpenStudy (anonymous):

I got as far as (1/cos^2x)+(sin^x/cos^2x) and the final answer is 2sec^2x-1. I need to know how to get between these steps specifically. Thanks in advanced.

OpenStudy (anonymous):

check this

OpenStudy (welshfella):

sec^2 x + tan^2x now sec^2 x = 1 + tan^2 x tan^2 x = sec^ 2x - 1 so we have sec^2 x+ sec^2 x - 1 = 2sec^2 x - 1